# OmniGen: Unified Image Generation

### v1

@@ 6 Limitations and Discussions
We believe that most limitations can be addressed by training the model on more related data. Moreover, compared to most models, fine-tuning OmniGen for downstream tasks is simpler, as it inherently supports various image generation tasks without the need for extensive efforts and costs to build additional networks.
7 Related Work
7.1 Generative Foundation Models
The generative foundation model serves as the core of many contemporary artificial intelligence systems, revolutionizing the way machines interact with humans. The GPT series [54; 48] have demonstrated that language models can learn numerous tasks via training on a large-scale dataset. Following this trend, the rise of large language models (LLMs) [43; 3; 1] has further showcased their versatility, adeptly performing various tasks such as question answering, text summarization, and code generation within a single framework. Beyond language, multimodal large language models [39; 12] have been proposed to integrate vision and language capabilities. For example, as a typical architecture and popular trend, LLaVA [39] equips the LLM with visual perception and understanding capabilities by linking the vision encoder to the LLM through a connector layer. These models have shown impressive performance in vision-language understanding tasks. However, despite their ability to handle mixed text and image inputs, they lack the capability to generate images. The construction of a universal foundation model for image generation remains unclear and has not been fully explored. In this work, we propose a universal generative model that accepts arbitrary interleaved multimodal inputs and generates images, marking a significant stride towards a general-purpose image generation foundation model.
Recently, some works have explored unified models that support both text and image generation. In Chameleon [61], images and texts are both tokenized into the token sequence and modeled via discrete autoregressive modeling. Concurrent works such as TransFusion [77] and Show-O [69] unify diffusion and autoregressive methods into a single model, generating text autoregressively and images through diffusion. Nonetheless, like most existing diffusion models, they can only perform text-to-image tasks and cannot handle more complex and various visual generation tasks. The unification of tasks in visual generation remains unexplored. Unlike these efforts, our current focus is on the unification of diverse visual generation tasks. The model is capable of performing various tasks, including text-to-image generation, image editing, subject-driven generation, virtual try-on, image deblurring, human pose recognition, and more. To the best of our knowledge, this is the first model capable of unifying such a wide range of visual generation tasks. Building on this foundation, further expansion into text generation is planned as the next step in the research agenda.
7.2 Diffusion Model
Recent advancements in diffusion models have been remarkable, with notable contributions from the Stable Diffusion series [56; 52; 13], DALL-E [55], and Imagen [26]. These models are predominantly designed for text-to-image generation tasks. To facilitate visual-conditioned generation, approaches such as ControlNet [73] and T2i-Adapter [45] introduce supplementary networks integrated into existing text-to-image models, thereby enabling them to accommodate image-based conditions. StyleShot [17] incorporates a style-sensitive encoder to manipulate the style feature of the output images. InstructPix2Pix [4] addresses image editing by augmenting the model with additional input channels. SEED-X [20] and Kosmos-G [49] employ an MLLM to replace the CLIP encoder in SD, improve the performance on the specific downstream task. However, these methods are task-specific, extending the capabilities of SD by modifying the model architecture. In contrast, OmniGen is a model that natively supports various image generative tasks, unifying all tasks into a single framework. Multi-task learning enhances the model’s capabilities and also leads to the emergence of new abilities. Furthermore, when addressing various real-world tasks, OmniGen no longer requires any preprocessing steps or assistance from other models.
There is some work exploring the unification of computer vision (CV) tasks [2; 65; 16; 21]. However, these efforts primarily focus on classic vision tasks and do not support general image generation tasks. Additionally, current models often underperform compared to those specifically designed and trained for corresponding tasks, limiting their practical applications in real-world scenarios. In our work, the introduction of CV tasks plays a crucial role in enabling the model to learn general knowledge, thereby enhancing its image-generation capabilities and fostering the emergence of new abilities. For instance, incorporating the human pose estimation task has led to the model’s ability to generate new images directly based on the pose of a reference image, without the need for an additional model to extract the human pose. At present, We are not obsessed with the pursuit of optimal scores on CV tasks, but leave the fine-tuning of CV task performance for future research.

### v2

@@ 9.5 Failure cases
We believe that most limitations can be addressed by training the model on more related data. Moreover, compared to most models, fine-tuning OmniGen for downstream tasks is simpler. Due to OmniGen’s inherent support for multimodal inputs, users do not need to spend significant effort on designing networks for specific tasks (for example, there is no need to add ControlNet for conditional generation, nor to design additional style encoders for image styles). We will open-source the training code, and users will only need to prepare their data to fine-tune it for various tasks, including new tasks that have never been supported by any model before. We believe this approach can inspire a wide range of interesting downstream applications.
Due to a lack of high-quality image data and insufficient model size, OmniGen’s image quality isn’t as good as the latest FLUX model. On one hand, expanding the data and model size is crucial; on the other hand, enabling existing models like FLUX to have similar multimodal instruction-following capabilities is also a potential research direction.
Figure 13: Qualitative results of image editing task. OmniGen achieves better results than other open-source models. Thanks to modeling both text and images within a unified transformer, OmniGen demonstrates strong multimodal instruction understanding and is capable of handling complex tasks, as shown in (f).
Figure 14: Qualitative results on DreamBench. OmniGen achieves better subject fidelity and text fidelity, preserving the object from the reference image while following new textual instructions. In contrast, Kosmos-G does not always follow text instructions, resulting in a higher CLIP-I similarity.
Figure 15: Results of Identity-Preserving Generation. OmniGen doesn’t require additional face detectors and encoders. OmniGen can utilize clothing information from input images, as shown in examples (a) and (b), and can also follow text instructions to generate new clothing, as seen in examples (c) and (d), demonstrating a higher degree of flexible control.
Figure 16: Some unique features of OmniGen. OmniGen can not only extract the desired object from image containing multiple objects based on instructions, but also process multiple images simultaneously. To our knowledge, no other model currently can offer this level of flexible and precise control.
Figure 17: Failure cases of OmniGen
Figure 18: Left: Illustration of the construction process for the GRIT-Entity dataset. We used instance segmentation and repainting methods to acquire a large volume of data. Right: Illustration of the cross-verification strategy used in constructing our web images dataset. For a group photo of Person-A and Person-B, we sampled several images from individual photos of Person-A and Person-B and asked MLLM whether they appear in the group photo. A group photo is retained only if the ”Yes” ratio for both Person-A and Person-B meets a specific threshold. The individual images marked as ”Yes” are then used to construct data pairs with the corresponding group image.
10 Subject-driven dataset
Figure 18-Left shows the process of constructing the GRIT-Entity dataset. Although the GRIT-based approach provides a substantial amount of data, the input data extracted directly from original images can lead the model to fall into simple copy-paste patterns. To fully unleash the subject-driven image generation capability of OmniGen, we constructed a high-quality web images training dataset using natural images of well-known individuals. First, we sampled 20 million Alt-text entries from the Datacomp dataset [^4] https://huggingface.co/datasets/UCSC-VLAA/Recap-DataComp-1B and used spaCy [^5] https://github.com/explosion/spaCy for named entity recognition. We selected the most frequently occurring names and employed GPT-4o to filter out real, notable individuals, resulting in 2,000 names. Furthermore, we expanded the initial 2,000 names by including closely related individuals, resulting in approximately 10,000 name pairs. We then scraped images of these individuals and pairs from search engines. Due to the noise in web images, where scraped images may not contain the specified individuals, we designed a cross-verification strategy using InternVL [10] to filter single and group images, as detailed in Figure 18-Right. The retained single and group images were then captioned with details such as attire and actions. We successfully build a dataset of 533K image pairs.